OpenStudy (anonymous):

a straight stretch of one lane highway is 1000 miles long. one truck starts at the south end and goes north at 40 mph. At the same time, another truck begins at the north end and does south at 60 mph. they eventually run into each other and stop. a fly flies at 80mph between the two trucks he begins on the front of one of the trucks,flies to the other truck. as soon as he gets to one truck, he turns around and flies to the other truck. When the trucks drives into each other the fly gets smashed . How far does the fly fly total?

OpenStudy (anonymous):

wow im stupid 80*10=800 miles

OpenStudy (anonymous):

y u say that and how you get that as a sum

OpenStudy (anonymous):

cuz it was an easier problem than i made it out to b. the trucks r 1000 miles apart right? 1 is traveling at 60 and 1 at 40 so together they travel at 100mi/hr

OpenStudy (anonymous):

thus they crash when they travel the 1000 miles together or t=1000/100

OpenStudy (anonymous):

so they crash in 10 hours

OpenStudy (anonymous):

ok cause your answer was correct i thought u didnt think it was

OpenStudy (anonymous):

now the fly has been traveling 80 miles per hr constantly. and it has 10 hours to do this so the amount of miles it travels is 80*10=800miles

OpenStudy (anonymous):

i was doing it a different way and then that answer popped in my head and made me feel bad